OpenStudy (anonymous):

What is the sum of the finite arithmetic series? (–5) + 0 + 5 + 10 + ... + 65

OpenStudy (raden):

use the formula : Sn = n/2 (a1 + an)

OpenStudy (anonymous):

i kinda dont understand how to use it

OpenStudy (raden):

but u have to find n, first

OpenStudy (anonymous):

can you show me how i would do that ?

OpenStudy (raden):

well, from the formula an = a1 + (n-1)d with a1 = the first term and d=difference given : an = 65, a1=-5, and its difference = 5 so a1 + (n-1)d = 65 -5 + (n-1)5 = 65 -5 + 5n - 5 = 65 5n - 10 = 65 5n = 65 + 10 5n = 75 n = 75/5 = 15

OpenStudy (raden):

it just get the n now, we can determine the sum of that series : Sn = n/2 (a1 + an) plug the values of n=15, a1=-5, an=65 to formula above S15 = 15/2 (-5 + 65) S15 = 15/2 (60) S15 = ...

OpenStudy (anonymous):

450?

OpenStudy (raden):

yes, that;s correct

OpenStudy (anonymous):

yayy thank youu soo much!! :)

OpenStudy (raden):

you're welcome